# Escaping Plato’s Cave: Towards the Alignment of 3D and Text Latent Spaces

### v1

@@ Abstract
Recent works have shown that, when trained at scale, uni-modal 2D vision and text encoders converge to learned features that share remarkable structural properties, despite arising from different representations. However, the role of 3D encoders with respect to other modalities remains unexplored. Furthermore, existing 3D foundation models that leverage large datasets are typically trained with explicit alignment objectives with respect to frozen encoders from other representations. In this work, we investigate the possibility of a posteriori alignment of representations obtained from uni-modal 3D encoders compared to text-based feature spaces. We show that naive post-training feature alignment of uni-modal text and 3D encoders results in limited performance. We then focus on extracting subspaces of the corresponding feature spaces and discover that by projecting learned representations onto well-chosen lower-dimensional subspaces the quality of alignment becomes significantly higher, leading to improved accuracy on matching and retrieval tasks. Our analysis further sheds light on the nature of these shared subspaces, which roughly separate between semantic and geometric data representations. Overall, ours is the first work that helps to establish a baseline for post-training alignment of 3D uni-modal and text feature spaces, and helps to highlight both the shared and unique properties of 3D data compared to other representations.
@@ 3.2 Alignment approaches
Affine + Subspace Projection PointBert 67.6 85.6 55.2 75.8 45.0 70.6
Affine + Subspace Projection SparseConv 67.6 87.0 57.2 75.6 46.8 69.2
@@
Table 1: Matching and retrieval performance across 3D and text encoders using different alignment approaches. We use 30,000 anchors for subspace projection and affine transformation approaches, and 1,000 anchors for local CKA. A query set of 500 is uniformly sampled, with results averaged over 3 different seeds. The subspace dimension is fixed at 50. SparseConv corresponds to MinkowskiNet. Our approach (Ours) consistently demonstrates improved matching and retrieval performance, with multi-modal 3D encoders setting the upper bound for performance. Additional top-$k$ retrieval metrics are provided in the supplementary.
@@ 4.2 Encoders
Multi-modal 3D Encoders.  For multi-modal pre-training, we use OpenShape [34] pre-trained models, trained on point cloud, image, and text triplets of Objaverse with a contrastive pre-text task to align 3D encoders with frozen CLIP encoders. OpenShape’s pipeline also includes enhanced captions, generated by applying Blip [31] on images and GPT-4 [1] for caption refinement. To evaluate the effect of different architectures, we include both a transformer-based model, PointBERT [67], and a sparse convolution-based model, MinkowskiNet [7], each representing distinct approaches in processing 3D data.
Uni-modal 3D Encoders.  For the uni-modal setup, we pre-train PointBERT on Objaverse using its original pretext tasks, which include masked point reconstruction and a uni-modal contrastive loss to encourage robust shape representation. We also explore two additional architectures: a MinkowskiNet model and a simpler architecture in PointNet [48], each pre-trained using a shape-level contrastive learning method [19]. This approach contrasts different partial views of input shapes. Across all 3D encoders in this setup, we fix the latent dimension at 512 to maintain consistency in representation space comparisons.
Text encoders.  We use the text encoder from OpenCLIP ViT-bigG-14 [25], chosen to match the multi-modal encoders used in OpenShape. Additionally, we examine alignment with purely uni-modal text encoders by including BERT [12] and RoBERTa [36], popular encoders that enable analysis of uni-modal text-3D representation alignment.
Across all pre-training setups, parameters are kept consistent to facilitate direct comparisons.
@@ 4.3 Downstream tasks
We evaluate our alignment approach on Objaverse-LVIS [18], a human-verified test subset of Objaverse that contains 1,156 object categories. This subset is specifically reserved for evaluation and is unseen during pre-training. The captions are generated with Cap3D [37], which provides enhanced descriptive text for each 3D shape. Our evaluation framework consists of two main tasks: matching and retrieval. The matching task involves finding the correct permutation of captions for perfect matching given a shuffled set of query images and their corresponding captions; we utilize the linear sum assignment approach to perform this task [30]. For the retrieval task, the model must identify the correct 3D object from the query set based on a text caption. These tasks are particularly effective in measuring the cross-modal capabilities of encoders, and have been evaluated in prior Vision-Text studies [39]. While our main results emphasize the matching and top-5 retrieval tasks, we provide comprehensive evaluation of top-1 and top-10 retrieval metrics in the supplementary.
Figure 2: Linear CKA scores between text and 3D encoders without alignment. SpConv corresponds to MinkowskiNet. Higher scores reflect stronger alignment between encoder pairs, with the strongest alignment observed between multi-modal 3D encoders and CLIP text encoder due to their shared training on aligned representations. Uni-modal 3D encoders show significantly lower alignment with text encoders, though slightly higher with the CLIP text encoder.
Figure 3: Linear CKA scores between text and 3D encoders after affine translation. SpConv refers to MinkowskiNet. Affine translation results in a similar increase in similarity for both 3D-to-text and text-to-3D directions.
@@ 5.2 Latent Space Alignment results
Existing Approaches Enable Limited Alignment  We first assess whether previous successful approaches—affine transformation and local CKA—can achieve meaningful 3D-text alignment. As shown in Tab. 1, both methods yield modest alignment improvements over the unaligned baseline, where unimodal 3D encoders start near zero in matching and retrieval tasks. These results suggest a small, yet notable, alignment shift. However, even with this improvement, alignment remains significantly lower than the alignment achieved in vision-text benchmarks [39]. This finding hints at the limits of uni-modal 3D encoders in achieving similar Vision-Text alignment performance, and might necessitate a different approach to possibly achieve notable alignment.
@@
Importance of subspace projection (Ours).  While aligning entire latent spaces of 3D and text encoders achieved some success, results remained well below vision-text alignment benchmarks. Motivated by this limitation, we propose aligning lower-dimensional subspaces, based on the hypothesis that 3D and text representations might intersect within a shared latent subspace. Using a validation set, we report in Fig. 4 the impact of subspace dimension, obtained through our CCA approach (see Sec. 3.3) combined with the affine transformation on the top-5 retrieval accuracy of uni-modal PointBert. Notably, while affine alignment alone yields better performance at higher dimensions, our subspace projection approach significantly outperforms it when the subspace dimension is reduced. This finding indicates that alignment quality is optimized within carefully chosen lower-dimensional spaces, reinforcing our hypothesis that 3D-text alignment is more successful within targeted subspaces. The results in Tab. 1 further validate this approach, showing a substantial increase in matching and retrieval performance across all 3D and text encoder pairs, outperforming both alignment approaches when they operate on the original latent space.
@@
3D encoder complexity’s low impact on alignment.  As anticipated, multi-modal 3D encoders reach the highest alignment performance across all metrics, effectively setting an upper bound. Interestingly, however, PointBERT outperforms the more complex SparseConv among uni-modal 3D encoders, suggesting that increasing model complexity alone does not guarantee improved alignment in 3D-text tasks. Surprisingly, even PointNet—a relatively simple model—achieves similar alignment scores, showing that factors other than model complexity may play a pivotal role in 3D-text alignment. This observation contrasts with vision-text alignment results [22, 39], where complex models typically leverage large datasets more effectively. Our findings thus indicate a distinctive aspect of 3D-text alignment: model simplicity does not impede, and may even aid, the interpretability and compatibility of learned features for cross-modal alignment.
@@
Figure 5: Effect of anchor set size on retrieval performance. Validation set results with a fixed subspace dimension $d=50$ show that retrieval performance improves as the anchor subset size increases but eventually reach a plateau. Results are shown for the uni-modal PointBERT and CLIP text encoder, with generalizations to other encoders provided in the supplementary.
@@
Increased geometric awareness of the text latent subspace.  To quantify the increase in geometric awareness within the projected text subspace, we calculate the Pearson correlation between the Chamfer distances of a query set of 500 shapes and the pairwise distances between feature vectors within the text subspace. We then compare these results to those obtained from the original text latent space. Results in Fig. 6 show that the optimal subspace dimension yields a notably high correlation with Chamfer distances, signaling a notable increase in the sensitivity of the text encoder to geometric properties when projected into this specific subspace.
@@ 6 Conclusion, Limitations and Future Work
Affine + Subspace Projection PointBert 56.4 90.8 38.8 81.8 32.4 78.8
Affine + Subspace Projection SparseConv 58.2 93.6 46.0 86.2 37.2 79.8
@@
Table 2: Top-1 and top-5 retrieval accuracy across 3D and text encoders using different alignment approaches. We use 30,000 anchors for subspace projection and affine transformation approaches, and 1,000 anchors for local CKA. A query set of 500 is uniformly sampled, with results averaged over 3 different seeds. The subspace dimension is fixed at 50. SparseConv corresponds to MinkowskiNet. Our approach (Ours) consistently demonstrates improved retrieval performance, with multi-modal 3D encoders setting the upper bound for performance.
This supplementary document provides additional details and results to complement the main material. Specifically, we first extend the matching and top-5 downstream results from Table 1 in the main paper by including top-1 and top-10 retrieval scores in App. A. Then, in App. B, we analyze the impact of chosen subspace dimensions and anchor counts across other pairs of text and 3D encoders. These additional results provide a more comprehensive understanding of our alignment approach and its performance across different setups.
Appendix A Downstream results
@@
Appendix B Additional ablations
@@
Number of anchors’ impact on alignment.  We find that aligning representations using our subspace projection approach generally requires a similar number of anchors across different pairs of text and 3D encoders, as we observe in Fig. 9. This consistency suggests that the alignment of latent spaces depends on a comparable amount of data, with limited scalability to very large datasets.
(a) CLIP and PointBert
(b) CLIP and MinkowskiNet
(c) CLIP and PointNet
(d) RoBERTa and PointBert
(e) RoBERTa and MinkowskiNet
(f) RoBERTa and PointNet
(g) BERT and PointBert
(h) BERT and MinkowskiNet
(i) BERT and PointNet
Figure 9: Effect of anchor set size on retrieval performance. Validation set results with a fixed subspace dimension $d=50$ show that retrieval performance improves as the anchor subset size increases but eventually reach a plateau. Each plot corresponds to a pair of Text Encoder and 3D Encoder. Anchors scalability is constant across different encoders settings.

### v2

@@ Abstract
Recent works have shown that, when trained at scale, uni-modal 2D vision and text encoders converge to learned features that share remarkable structural properties, despite arising from different representations. However, the role of 3D encoders with respect to other modalities remains unexplored. Furthermore, existing 3D foundation models that leverage large datasets are typically trained with explicit alignment objectives with respect to frozen encoders from other representations. In this work, we investigate the possibility of a posteriori alignment of representations obtained from uni-modal 3D encoders compared to text-based feature spaces. We show that naive post-training feature alignment of uni-modal text and 3D encoders results in limited performance. We then focus on extracting subspaces of the corresponding feature spaces and discover that by projecting learned representations onto well-chosen lower-dimensional subspaces the quality of alignment becomes significantly higher, leading to improved accuracy on matching and retrieval tasks. Our analysis further sheds light on the nature of these shared subspaces, which roughly separate between semantic and geometric data representations. Overall, ours is the first work that helps to establish a baseline for post-training alignment of 3D uni-modal and text feature spaces, and helps to highlight both the shared and unique properties of 3D data compared to other representations. Our code and weights are available at https://github.com/Souhail-01/3d-text-alignment.
@@ 3.2 Alignment approaches
Affine + Subspace Projection OpenShape 67.6 85.6 55.2 75.8 45.0 70.6
Affine + Subspace Projection ULIP-2 65.6 85.2 47.2 70.6 35.2 59.8
Affine + Subspace Projection Uni3D 61.4 81.6 45.8 63.8 34.4 47.2
@@
Table 1: Matching and retrieval performance across 3D and text encoders using different alignment approaches. We use 30,000 anchors for subspace projection and affine transformation approaches, and 1,000 anchors for local CKA. A query set of 500 is uniformly sampled, with results averaged over 3 different seeds. The subspace dimension is fixed at 50. Our approach (Ours) consistently demonstrates improved matching and retrieval performance, with multi-modal 3D encoders setting the upper bound for performance. Additional top-$k$ retrieval metrics are provided in the supplementary.
@@ 4.2 Encoders
Multi-modal 3D Encoders.  For multi-modal pre-training, we use OpenShape, ULIP-2 and Uni3D [32] pre-trained models, trained on point cloud, image, and text triplets of Objaverse with a contrastive pre-text task to align 3D encoders with frozen CLIP encoders. We adopt the Point-BERT-based variant of each model [66]. We primarily focus on OpenShape for simplicity but generalize results to ULIP-2 and Uni3D.
Uni-modal 3D Encoders.  For the uni-modal setup, we pre-train PointBERT on Objaverse using its original pretext tasks, which include masked point reconstruction and a uni-modal contrastive loss to encourage robust shape representation. We also explore two additional architectures: a Sparse convolution (MinkowskiNet [6]) model and a simpler architecture in PointNet [46], each pre-trained using a shape-level contrastive learning method [18]. This approach contrasts different partial views of input shapes. Across all 3D encoders in this setup, we fix the latent dimension at 512 to maintain consistency in representation space comparisons.
Text encoders.  We use the text encoder from OpenCLIP ViT-bigG-14 [24], chosen to match text encoder used in OpenShape. Additionally, we examine alignment with purely uni-modal text encoders by including BERT [11] and RoBERTa [34], we also evaluate the alignement with T5 [48] in the supplementary.
Across all pre-training setups, parameters are kept consistent to facilitate direct comparisons. We detail the technical details in the supplementary.
@@ 4.3 Downstream tasks
We evaluate our alignment approach on Objaverse-LVIS [17], a human-verified test subset of Objaverse that contains 1,156 object categories. This subset is specifically reserved for evaluation and is unseen during pre-training. The captions are generated with Cap3D [35], which provides enhanced descriptive text for each 3D shape. Our evaluation framework consists of two main tasks: matching and retrieval. The matching task involves finding the correct permutation of captions for perfect matching given a shuffled set of query images and their corresponding captions; we utilize the linear sum assignment approach to perform this task [29]. For the retrieval task, the model must identify the correct 3D object from the query set based on a text caption. These tasks are particularly effective in measuring the cross-modal capabilities of encoders, and have been evaluated in prior Vision-Text studies [37]. While our main results emphasize the matching and top-5 retrieval tasks, we provide evaluation of top-1 and top-10 retrieval metrics in the supplementary.
Figure 2: Linear CKA scores between text and 3D encoders without alignment. Higher scores reflect stronger alignment between encoder pairs, with the strongest alignment observed between multi-modal 3D encoders (OpenShape) and CLIP text encoder due to their shared training on aligned representations. Uni-modal 3D encoders show significantly lower alignment with text encoders, although slightly higher with the CLIP text encoder.
Figure 3: Linear CKA scores between text and 3D encoders after affine translation. Affine translation results in a consistent increase in similarity for both 3D-to-text and text-to-3D directions.
@@ 5.2 Latent Space Alignment results
Existing approaches enable limited alignment  We first assess whether previous successful approaches—affine transformation and local CKA—can achieve meaningful 3D-text alignment. As shown in Tab. 1, both methods yield modest alignment improvements over the unaligned baseline, where uni-modal 3D encoders start near zero in matching and retrieval tasks. These results suggest a small alignment shift. However, even with this improvement, alignment remains significantly lower than the alignment achieved in vision-text benchmarks [37]. This finding hints at the limits of uni-modal 3D encoders in achieving similar Vision-Text alignment performance, and might necessitate a different approach to align their latent space with text encoders.
@@
Importance of subspace projection (Ours).  While aligning the latent spaces of 3D and text encoders achieved some success with existing methods, results remained well below vision-text alignment benchmarks. Motivated by this limitation, we propose aligning lower-dimensional subspaces, based on the hypothesis that 3D and text representations might intersect within a shared latent subspace. Using a validation set, we report in Fig. 4 the impact of subspace dimension, obtained through our CCA approach (see Sec. 3.3) combined with the affine transformation on the top-5 retrieval accuracy of uni-modal PointBert. While affine alignment alone yields better performance at higher dimensions, our subspace projection approach significantly outperforms it when the subspace dimension is reduced. This finding indicates that alignment quality is optimized within carefully chosen lower-dimensional spaces, reinforcing our hypothesis that 3D-text alignment is more successful within targeted subspaces. The results in Tab. 1 further validate this approach, showing a substantial increase in matching and retrieval performance across all 3D and text encoder pairs, outperforming both alignment approaches when they operate on the original latent space.
Multi-modal encoders as an upper bound of performance.  We include 3D multi-modal encoders for two main reasons: (1) As an upper bound for alignment performance, (2) To study how alignment degrades across different text encoders, including those that were not used during pre-training. In this context, the inclusion of ’Affine + Subspace Projection’ multi-modal results in Tab. 1 shows how using our approach significantly narrows the gap between multi-modal and uni-modal performance. However, to illustrate point 1. above, we also measured with cosine similarity the alignment of OpenShape (a multi-modal 3D encoder) and the same CLIP text encoder used during pre-training, achieving 0.94 for top-5 retrieval accuracy, an expected result since 3D and text encoders are explictly aligned during pre-training.
@@
3D encoder complexity’s low impact on alignment.  PointBERT outperforms the more complex SparseConv among uni-modal 3D encoders, suggesting that increasing model complexity alone does not guarantee improved alignment in 3D-text tasks. Surprisingly, even PointNet—a relatively simple model—achieves similar alignment scores, showing that factors other than model complexity may play a pivotal role in 3D-text alignment. This observation contrasts with vision-text alignment results [21, 37], where complex models typically leverage large datasets more effectively. Our findings thus indicate a distinctive aspect of 3D-text alignment: model simplicity does not impede, and may even aid, the interpretability and compatibility of learned features for cross-modal alignment.
@@
Figure 5: Effect of anchor set size on retrieval performance. Validation set results with different subspace dimensions show that retrieval performance improves as the anchor subset size increases but eventually reaches a plateau. Results are shown for the uni-modal PointBERT and CLIP text encoder.
@@
Increased geometric awareness of the text latent subspace.  To quantify the increase in geometric awareness within the projected text subspace, we compute the Pearson correlation between the Chamfer distances of a query set of 500 shapes and the pairwise distances between feature vectors within the text subspace and compare these results to those obtained from the original text latent space. Results in Fig. 6 show that the optimal subspace dimension yields a high correlation with Chamfer distances, signaling an increase in the sensitivity of the text encoder to geometric properties when projected into this specific subspace.
@@ 6 Conclusion, Limitations and Future Work
Affine + Subspace Projection OpenShape 56.4 90.8 38.8 81.8 32.4 78.8
Affine + Subspace Projection ULIP-2 54.2 90.2 37.0 80.8 29.2 69.2
Affine + Subspace Projection Uni3D 47.0 89.0 29.4 73.8 19.0 59.4
@@
Table 2: Top-1 and top-5 retrieval accuracy across 3D and text encoders using different alignment approaches. We use 30,000 anchors for subspace projection and affine transformation approaches, and 1,000 anchors for local CKA. A query set of 500 is uniformly sampled, with results averaged over 3 different seeds. The subspace dimension is fixed at 50. Our approach (Ours) consistently demonstrates improved retrieval performance, with multi-modal 3D encoders setting the upper bound for performance.
Method 3D Encoder T5
Matching accuracy Top-5 retrieval
Multi-modal 3D encoder
Affine + Subspace Projection (Ours) OpenShape 65.0 82.6
Affine + Subspace Projection (Ours) ULIP-2 51.8 73.2
Affine + Subspace Projection (Ours) Uni3D 53.6 67.0
Uni-modal 3D encoder
Affine + Subspace Projection (Ours) PointBert 21.6 28.4
Affine + Subspace Projection (Ours) SparseConv 22.8 23.2
Affine + Subspace Projection (Ours) Pointnet 21.6 22.0
Table 3: Matching and Top-5 retrieval accuracy using T5 text encoder and different 3D encoders. The Affine + Subspace Projection (Ours) method is evaluated across both multi-modal and uni-modal 3D encoders. T5 is better aligned to 3D encoders compared to other uni-modal text encoders as presented in Table 1 of the main paper.
This supplementary document provides additional technical details and experimental results to complement the main material. Specifically, we first outline some technical details pertinent to our framework in App. A. Then, we evaluate in App. B the alignment performance using a different text encoder, T5. Next, we extend the matching and top-5 downstream results from Table 1 in the main paper by including top-1 and top-10 retrieval scores in App. C. Furthermore, in App. D, we examine the impact of chosen subspace dimensions across different pairs of text and 3D encoders. These additional results provide a more comprehensive understanding of our alignment approach and its robustness across different configurations.
Appendix A Implementation Details
In our framework, we use mean pooling for text encoders to obtain a fixed-size text representation. While we also experimented with using a class token for text encoding, it yielded consistently similar results. For 3D encoders, we extract the global output feature as the final representation.
The embedding size is set to 512 whenever possible, with the option to use projection layers when necessary. However, for multi-modal pre-trained models and certain text encoders, the embedding dimension may vary. In such cases, we adopt two distinct strategies: (1) for Canonical Correlation Analysis (CCA)-based approaches (Ours), the maximum subspace dimension is determined as the minimum of the embedding sizes of the 3D and text encoders; (2) for affine transformation-based alignment, we follow prior work by padding the lower-dimensional representation to match the higher-dimensional one.
For training parameters and dataset configurations related to uni-modal encoders, we adhere to the OpenShape settings, ensuring consistency with existing benchmarks.
Appendix B Evaluation with an Additional Text Encoder
To further evaluate the generalization of our alignment approach, we test it using an additional text encoder, T5. Unlike BERT, RoBERTa, and CLIP’s text encoder, which are encoder-only architectures, T5 follows an encoder-decoder structure. We average its encoder output embeddings and use the resulting vector as the text representation.
As shown in Tab. 3, while T5 does not achieve the same performance as CLIP’s text encoder, it consistently outperforms other uni-modal text encoders, such as BERT and RoBERTa, in alignment tasks. These results suggest that the encoder-decoder structure may provide richer text representations for cross-modal alignment, although multi-modal encoders like CLIP text encoder remain superior for this task.
Appendix C Downstream Results
@@
The alignment approaches studied thus far exhibit limited generalization to the zero-shot classification downstream task. In particular, top-1 accuracy on Objaverse-LVIS remains below 3% even for the best-performing uni-modal 3D encoders when aligned with text encoders which is way lower to the 40% and more attained by OpenShape. This performance bottleneck can be attributed primarily to the repetitive nature of the captions used in this task: instances within the same class often share identical or nearly identical textual descriptions, leading to duplicated text embeddings. This opens up a new direction to enhance these approaches with zero-shot classification capabilities.
@@
Appendix D Additional Ablations
@@
Experiments Favoring Uni-Modal Encoders:  Although our focus is on bridging the gap between uni-modal latent spaces, we can demonstrate that uni-modal 3D encoders have an edge over multi-modal encoders in geometry understanding. For instance, the Pearson correlation experiment (Fig. 9 which complements Fig. 6 in the main) shows that the subspaces emerging from text 3D multi-modal encoders do not correlate with geometric similarity, thus potentially limiting their utility in tasks, where geometric similarity is important.
Figure 9: Pearson correlation between shape query Chamfer distances and pairwise distances. We observe a higher Pearson correlation in the projected text uni-modal latent subspace compared to the projected text multi-modal subspace.